Monopolistically firm sells balls. What will be the profit?



Price Data quantity Data output Total cost


$11 6 6 $48


$10 7 7 $52


$9 8 8 $57


$8 9 9 $63


$7 10 10 $70




1
Expert's answer
2021-12-03T11:08:46-0500

P Q TR MR TC MC

$11 6 $66 - $48 -

$10 7 $70 4 $52 4

$9 8 $72 2 $57 5

$8 9 $72 0 $63 6

$7 10 $70 -2 $70 7

The profit-maximizing output occurs when MR = MC = $4, so Q = 7, P = $10, TP = TR - TC = 70 - 52 = $18.